# Empowering Agentic Video Analytics Systems with Video Language Models

### v3

@@ 7.3.5. System Overhead
Avas is designed to enable the near real-time construction of EKGs. As shown in Fig. 10, we measured the average processing speed (in FPS) of Avas while constructing EKGs from LVBench videos across various hardware platforms, with the input video stream fixed at 2 FPS. On 2 $\times$ A100 GPUs, Avas achieved an impressive processing speed of 6.7 FPS, significantly exceeding the input stream rate. On a single RTX 4090, a typical edge server hardware, Avas maintained a processing speed of 4.4 FPS, still well above the input frame rate. Even on a single RTX 3090, Avas performed effectively, achieving 2.5 FPS. This performance demonstrates its capability to support efficient, near real-time EKG construction for L4 video analytics. The overhead during the retrieval and generation phases will be discussed further in §7.4.
To evaluate the generation overhead of Avas, we performed a latency breakdown of the key stages in the generation pipeline, as detailed in Tab. 2. The results indicate that a single retrieval operation is very swift, requiring only 0.44 seconds. This efficiency stems from the fact that this step involves vector matching between information established during the graph construction phase. The tree search stage accounts for a substantial portion of the overall generation latency. When employing Qwen2.5-14B, this stage exhibits a latency of 101.5 seconds. This duration increases correspondingly to 174.2 seconds when switching to the larger Qwen2.5-32B, which has a greater number of parameters. For check and answer (CA) stage, Qwen2.5-VL-7B demostrates a latency of 85.2 seconds, while the API-based Gemini-1.5-Pro shows a latency of 50.2 seconds.
@@
Figure 10. Total index construction overhead evaluated on various types of typical edge server hardware.
Stage Model Latency(s)
Retrieval JinaCLIP 0.44
Tree Search Qwen2.5-14B 101.5
Qwen2.5-32B 174.2
Check and Answer Qwen2.5-VL-7B 85.2
Gemini-1.5-Pro 50.2
Table 2. The latency breakdown for the retrieval and generation phase of Avas.
7.4. Ablation Evaluation
@@ 7.4.1. Different Index Construction Methods
Avas 39.7 0.31
Table 3. The achieved accuracy and construction overhead evaluated when using EKG and KG as index in Avas and baseline models on the subset of LVBench. The total video duration is around 1.2 hours.
We compare Avas’s Event Knowledge Graph (EKG) construction method with two representative knowledge graph-based construction methods: LightRAG (Guo et al., 2024) and MiniRAG (Fan et al., 2025). Since both of them only support text-only construction, we use the full set of descriptions obtained through the semantic chunking (§4.2) as their input textual corpus. We use Qwen2.5 7B to construct EKG and KG for Avas and baselines, respectively. In the retrieval and generation phase, we use the same LLM, Qwen2.5 14B and the same settings, e.g., maximum number of tokens of retrieved events or entities.
As shown in Table 3, Avas demonstrates a significant performance advantage over the baselines. Specifically, it achieves 11.6% higher accuracy than MiniRAG and 9.1% higher accuracy than LightRAG. Crucially, this improved performance comes with substantially less construction overhead, requiring only 0.31 hours compared to 3.49 and 3.52 hours for the baselines. The rationale is that baselines construct KG based on massive uniform chunks, while Avas utilizes the semantic chunks. This substantial gap in both effectiveness and efficiency highlights that Avas’s EKG construction method not only yields higher-quality knowledge representations but also drastically reduces the time needed to build the graph.
7.4.2. Different Tree Search Depths
@@
Avas(Qwen2.5 14B) 34.1 36.1 40.9 39.5
Avas(Qwen2.5 14B + Qwen2.5VL 7B) 49.3 52.1 53.8 50.2
Avas(Qwen2.5 14B + Gemini-1.5-Pro) 54.2 58.4 61.5 52.7
@@
Table 4. The achieved accuracy and overhead when applying different tree search depths in the agentic search of Avas evalauted on the subset of LVBench.
We also evaluate the achieved accuracy and overhead applying different settings in the retrieval phase, i.e., tree search depth. The effect of tree depth lies in a crucial trade-off: while shallower depths may struggle to retrieve comprehensive information, increasing the depth allows access to richer information from deeper nodes. However, this comes with a significant increase in tree search overhead, and the information from deeper levels can introduce more noise, potentially negatively impacting the final generation quality. Table 4 presents the results comparing different tree search depths on performance and tree search overhead. As shown, performance generally increases with increasing tree depth up to a certain point. Specifically, for all three Avas configurations evaluated, the highest accuracy is achieved at a tree search depth of 3. Accuracy decreases when the depth is further increased to 4, suggesting that excessive depth leads to the retrieval of detrimental noise or irrelevant information, outweighing the benefit of additional context. Conversely, the tree search overhead increases sharply with depth. Expanding the search from depth 1 (6.7s) to depth 2 (27.3s) incurs a moderate increase. Comparing the accuracy improvements and the overhead increase, a tree search depth of 3 offers the optimal balance.
7.4.3. Different Consistency Evaluation Settings
@@
For consistency-enhanced generation, Avas incorporates two key parameters: $\alpha$, which governs the balance between the contributions of thought consistency and answer consistency, and the number of generations for self-consistency evaluation. Fig. 11(a) illustrates the impact of varying $\alpha$ values on the accuracy achieved by Avas. Notably, the optimal performance is observed when $\alpha$ is set to $0.3$, highlighting the importance of jointly considering both intermediate thought consistency and final answer consistency to ensure robust results. As depicted in Fig. 11(b), the accuracy of Avas gradually improves as the number of self-consistency iterations increases. However, this improvement comes at the expense of significantly higher computational overhead. For example, increasing the self-consistency iterations from 8 to 16 yields only a 0.9% accuracy gain, while nearly doubling the computational cost. This demonstrates a clear trade-off between marginal accuracy improvements and resource efficiency. Balancing this trade-off, we adopt 8 self-consistency iterations in the implementation of Avas, ensuring a practical balance between performance and computational overhead.
Figure 11. The performance of Avas under varying consistency evaluation settings on the subset of LVBench.
8. Limitations and Future Work
@@
There are also limitations in the current design of Avas, which we explore for future work. Specifically: 1) The existing agentic retrieval and generation mechanism relies on a fixed tree-search strategy based on the Monte Carlo approach. While effective, this method is computationally expensive. The trajectories collected during the search process could be leveraged as training data to develop a model capable of dynamically selecting optimal search actions and depths based on the query and context. 2) Although the integrated VLM demonstrates robust general video understanding and reasoning capabilities, it may encounter challenges in certain specialized visual tasks, such as precise object counting. Incorporating lightweight, task-specific vision models as tools within the system could improve accuracy for such queries. Our future work will focus on enabling the VLM, functioning as an autonomous agent, to intelligently invoke these specialized tools, thereby addressing its limitations in handling specific tasks.
9. Conclusion
@@
This paper presents Avas, an advanced L4 video analytics system powered by VLMs. Avas enables comprehensive understanding and open-ended query analysis of large-scale, long-duration video data, overcoming the constraints of existing video analytics systems that are predominantly tailored to specific, pre-defined tasks. The system introduces novel designs, including near-real-time Event Knowledge Graph (EKG) index construction and an agentic retrieval and generation mechanism, facilitating efficient organization and analysis of extended video content to address complex queries. We demonstrate Avas’s superior performance on public video understanding benchmarks, as well as on our newly proposed benchmark, Avas-100, specifically designed to evaluate video analytics tasks.

### v5

@@ 7.3.5 System Overhead
Ava is designed to enable the near real-time construction of EKGs. As shown in Fig. 11, we measured the average processing speed (in FPS) of Ava while constructing EKGs from LVBench videos across various hardware platforms, with the input video stream fixed at 2 FPS. On 2 $\times$ A100 GPUs, Ava achieved an impressive processing speed of 6.7 FPS, significantly exceeding the input stream rate. On a single RTX 4090, a typical edge server hardware, Ava maintained a processing speed of 4.4 FPS, still well above the input frame rate. Even on a single RTX 3090, Ava performed effectively, achieving 2.5 FPS. This performance demonstrates its capability to support efficient, near real-time EKG construction for L4 video analytics. The overhead during the retrieval and generation phases will be discussed further in §7.4.
@@
Figure 11: Total index construction overhead evaluated on various types of typical edge server hardware.
To evaluate system overhead during the generation phase, we conducted a detailed latency analysis of three stages within the generation pipeline, as summarized in Table 2. The experiment is performed on one A100 GPU with LMDeploy [10] [^2] We configure cache_max_entry_countcache\_max\_entry\_count to 0.3 in LMDeploy, restricting the KV cache memory consumption to no more than 30% of the total GPU memory. The retrieval stage utilizing JinaCLIP demonstrates notable efficiency, requiring only 0.44 seconds per query and consuming less than 1 GB of GPU memory on average. The agentic search stage presents the predominant source of latency: 101.5 seconds with Qwen2.5-14B, rising to 174.2 seconds with the larger Qwen2.5-32B, thereby underscoring the inference time of LLMs as the principal runtime overhead. During the consistency-enhanced generation stage, Qwen2.5-VL-7B incurs a latency of 45.8 seconds and approximately 31 GB of memory usage, whereas the API-based Gemini-1.5-Pro achieves substantially lower latency at 14.2 seconds. The results demonstrate that the agentic searching stage represents the principal performance bottleneck, underscoring the importance of strategic model selection to balance accuracy and computational efficiency. Furthermore, the results also highlight the imperative of optimizing the agentic searching process, which we identify as a key direction for future research.
Stage Model Latency (s) GPU Memory (GB)
Tri-View Retrieval JinaCLIP 0.44 0.8
Agentic Searching Qwen2.5-14B 101.5 30
Qwen2.5-32B 174.2 40
Consistency Enhanced Gen. Qwen2.5-VL-7B 45.8 31
Gemini-1.5-Pro 14.2 -
Table 2: The breakdown of latency and GPU memory consumption across the three stages of the generation phase in Ava. All measurements are performed on a single A100 GPU, with LLMs deployed via LMDeploy [10] utilizing AWQ [24].
7.4 Ablation Evaluation
@@
Ava 39.7 0.31
Table 3: The achieved accuracy and construction overhead evaluated when using EKG and KG as index in Ava and baseline models on the subset of LVBench. The total video duration is around 1.2 hours.
7.4.1 Different Index Construction Methods
We compare Ava’s EKG construction method with two representative knowledge graph-based construction methods: LightRAG [16] and MiniRAG [13]. Since both of them only support text-only construction, we use the full set of descriptions obtained through the semantic chunking (§4.2) as their input textual corpus. We use Qwen2.5 7B to construct EKG and KG for Ava and baselines, respectively. In the retrieval and generation phase, we use the same LLM, Qwen2.5 14B and the same settings, e.g., maximum number of tokens of retrieved events or entities.
As shown in Table 3, Ava demonstrates a significant performance advantage over the baselines. Specifically, it achieves 11.6% higher accuracy than MiniRAG and 9.1% higher accuracy than LightRAG. Crucially, this improved performance comes with substantially less construction overhead, requiring only 0.31 hours compared to 3.49 and 3.52 hours for the baselines. The rationale is that baselines construct KG based on massive uniform chunks, while Ava utilizes the semantic chunks. This substantial gap in both effectiveness and efficiency highlights that Ava’s EKG construction method not only yields higher-quality knowledge representations but also drastically reduces the time needed to build the graph.
7.4.2 Different Tree Search Depths
We also evaluate the achieved accuracy and overhead applying different settings in the retrieval phase, i.e., tree search depth. The effect of tree depth lies in a crucial trade-off: while shallower depths may struggle to retrieve comprehensive information, increasing the depth allows access to richer information from deeper nodes. However, this comes with a significant increase in tree search overhead, and the information from deeper levels can introduce more noise, potentially negatively impacting the final generation quality. Table 4 presents the results comparing different tree search depths on performance and tree search overhead. As shown, performance generally increases with increasing tree depth up to a certain point. Specifically, for all three Ava configurations evaluated, the highest accuracy is achieved at a tree search depth of 3. Accuracy decreases when the depth is further increased to 4, suggesting that excessive depth leads to the retrieval of detrimental noise or irrelevant information, outweighing the benefit of additional context. Conversely, the tree search overhead increases sharply with depth. Expanding the search from depth 1 (6.7s) to depth 2 (27.3s) incurs a moderate increase. Comparing the accuracy improvements and the overhead increase, a tree search depth of 3 offers the optimal balance.
@@ 7.4.2 Different Tree Search Depths
Ava(Qwen2.5 14B) 34.1 36.1 40.9 39.5
Ava(Qwen2.5 14B + Qwen2.5VL 7B) 49.3 52.1 53.8 50.2
Ava(Qwen2.5 14B + Gemini-1.5-Pro) 54.2 58.4 61.5 52.7
@@
Table 4: The achieved accuracy and overhead when applying different tree search depths in the agentic search of Ava evaluated on the subset of LVBench.
7.4.3 Different Consistency Evaluation Settings
@@
For consistency-enhanced generation, Ava incorporates two key parameters: $\lambda$, which governs the balance between the contributions of thought consistency and answer consistency, and the number of generations for self-consistency evaluation. Fig. 12(a) illustrates the impact of varying $\lambda$ values on the accuracy achieved by Ava. Notably, the optimal performance is observed when $\lambda$ is set to $0.3$, highlighting the importance of jointly considering both intermediate thought consistency and final answer consistency to ensure robust results. As depicted in Fig. 12(b), the accuracy of Ava gradually improves as the number of self-consistency iterations increases. However, this improvement comes at the expense of significantly higher computational overhead. For example, increasing the self-consistency iterations from 8 to 16 yields only a 0.9% accuracy gain, while nearly doubling the computational cost. This demonstrates a clear trade-off between marginal accuracy improvements and resource efficiency. Balancing this trade-off, we adopt 8 self-consistency iterations in the implementation of Ava, ensuring a practical balance between performance and computational overhead.
8 Limitations and Future Work
@@
There are also limitations in the current design of Ava, which we explore for future work. Specifically: 1) The existing agentic retrieval and generation mechanism relies on a fixed tree-search strategy based on the Monte Carlo approach. While effective, this method is computationally expensive. The trajectories collected during the search process could be leveraged as training data to develop a model capable of dynamically selecting optimal search actions and depths based on the query and context. 2) Although the integrated VLM demonstrates robust general video understanding and reasoning capabilities, it may encounter challenges in certain specialized visual tasks, such as precise object counting. Incorporating lightweight, task-specific vision models as tools within the system could improve accuracy for such queries. Our future work will focus on enabling the VLM, functioning as an autonomous agent, to intelligently invoke these specialized tools, thereby addressing its limitations in handling specific tasks.
(a) Balance between thoughts and answer consistency
(b) Trade-offs using different self-consistency times
Figure 12: Performance of Ava under varying consistency evaluation settings on a subset of LVBench.
9 Conclusion
@@
This paper presents Ava, an advanced L4 video analytics system powered by VLMs. Ava enables comprehensive understanding and open-ended query analysis of large-scale, long-duration video data, overcoming the constraints of existing video analytics systems that are predominantly tailored to specific, pre-defined tasks. The system introduces novel designs, including near-real-time Event Knowledge Graph index construction and an agentic retrieval and generation mechanism, facilitating efficient organization and analysis of extended video content to address complex queries. We demonstrate Ava’s superior performance on public video understanding benchmarks, as well as on our newly proposed benchmark, Ava-100, specifically designed to evaluate video analytics tasks.
Acknowledgements
This work is partially supported by NSFC under grant No. 92467301, Key Research and Development Program of Zhejiang Province (Grant No: 2025C01012), and the ZJUCSE-Enflame cloud and edge intelligence joint laboratory.
Appendix A Ava-100 Benchmark
A.1 Benchmark overview
Figure 13: Scenarios and QA examples in Ava-100.
Ava-100 is established as a benchmark to advance research in L4 video analytics systems, with a particular focus on ultra-long-duration video understanding and complex reasoning tasks. In contrast to existing datasets that primarily focus on short video clips or domain-specific activities, Ava-100 centers on real-world, continuous video streams often spanning several hours. This design enables the evaluation of video analytics systems under realistic and challenging scenarios, including long-horizon temporal reasoning, sparse event detection, and cross-segment summarization, etc.
The benchmark currently comprises eight ultra-long videos encompassing diverse scenarios including egocentric recordings of daily activities, city tours, as well as third-person fixed-camera capturing traffic monitoring and wildlife monitoring. Each video is annotated with multiple question–answer (QA) pairs: human annotators generate ground-truth QAs, while GPT-4o is employed to produce multiple-choice options, subsequently verified through manual review. In total, Ava-100 contains approximately 100 hours of video content and 120 annotated QA pairs, offering a valuable resource for the evaluation of both video-language models and specialized video analytics systems.
Fig. 13 illustrates the representative scenarios and QA examples featured in Ava-100. Table 5 presents detailed statistics of the Ava-100 dataset, encompassing video duration and the number of QA pairs. The benchmark is publicly available at https://huggingface.co/datasets/iesc/Ava-100.
A.2 Selected Scenarios and Data Sources
A.2.1 Human Daily Activities
The egocentric daily-life scenario videos in Ava-100 are primarily sourced from the Ego4D dataset [15]. Ego4D is one of the largest existing egocentric video datasets, containing more than 3,600 hours of video collected across multiple countries, focusing on a wide range of daily activities such as household routines, cooking, shopping, and social interactions. It was originally designed to support research in egocentric perception, including tasks like action recognition, episodic memory, forecasting, and hand–object interaction understanding.
Video ID Duration (hours) #QA Pairs Views
ego-1 12.7 22 First-person (moving)
ego-2 11.7 19 First-person (moving)
citytour-1 12.0 19 First-person (moving)
citytour-2 10.5 20 First-person (moving)
traffic-1 14.9 12 Third-person (fixed)
traffic-2 13.9 13 Third-person (fixed)
wildlife-1 12.0 8 Third-person (fixed)
wildlife-2 11.5 7 Third-person (fixed)
Total 99.2 120
Table 5: Statistics of the Ava-100 dataset, including video duration, number of QA pairs, and perspective attributes for each video. The first four entries correspond to egocentric videos capturing moving perspectives, whereas the last four represent fixed third-person recordings.
Ego4D does not typically provide single, continuous clips exceeding 10 hours in duration. To construct long-form egocentric videos for Ava-100, we employed the concatenation strategy. Specifically, we prioritized the preservation of event diversity over maintaining the identity of the camera wearer. Accordingly, we selected and concatenated sub-clips from distinct Ego4D videos to generate extended sequences surpassing 10 hours. Throughout this process, we guarantee that the concatenated sub-clips encompassed a wide spectrum of human activities and contexts, thereby enhancing the richness of the video content while mitigating potential annotation ambiguities.
A.2.2 City Walking
The first-person city tour videos in Ava-100 are sourced from the YouTube channel 4K World Wandering[1], which features high-resolution walking tour videos recorded in various locations around the globe. These videos capture natural human navigation experiences in public spaces, making them a valuable complement to the daily-life egocentric content. The videos typically span up to four hours in length. To construct long-form sequences for Ava-100, we adopted the same concatenation strategy as described above. Specifically, we aimed to compose extended videos by combining multiple sub-clips, while preserving both geographic and contextual coherence.
In particular, we curated sub-videos predominantly from the same region or country to construct each long-form video. Meanwhile we intentionally enhanced diversity by integrating segments filmed under various weather conditions (e.g., sunny, snowy, rainy) and at different times of day (e.g., daytime, night) into the concatenated sequences. This design enriches the video contents, rendering the city walking videos in Ava-100 more representative of real-world navigation scenarios across diverse geographic and temporal conditions.
The city walking videos prioritize large-scale spatial cognition, long-term trajectory tracking, and environmental variability. Consequently, they present added challenges for QA annotation, as the questions frequently require reasoning about landmarks, navigation routes, or temporal dynamics within the environment.
A.2.3 Traffic Monitoring
The traffic monitoring videos in Ava-100 are sourced from the Bellevue Traffic dataset [7], which comprises continuous, fixed-camera recordings of real-world traffic scenes. This dataset is particularly well-suited for the study of long-term temporal dynamics, vehicle-pedestrian interactions, and congestion analysis.
For Ava-100, we selected two intersections: Bellevue_150th_Newport (spanning from 2017-09-11 03:08:29 to 2017-09-11 16:08:32) and Bellevue_Bellevue_NE8th (spanning from 2017-09-11 03:08:29 to 2017-09-11 15:08:32). These extensive, continuous recordings encompass a broad spectrum of traffic events across different periods of the day, including both peak hours and quieter intervals, thereby providing a rich and diverse dataset.
To construct QA pairs, we target the fine-grained traffic events, such as vehicle movements, pedestrian crossings, and congestion incidents. For example, whether a red car passed through the intersection between 4:30 and 4:40 and in which direction it traveled; the presence of pedestrians or cyclists at specific times. Such temporally anchored, detail-oriented QAs highlight the unique value of traffic monitoring videos for evaluating long-duration reasoning.
A.2.4 Wildlife Monitoring
The wildlife monitoring videos in Ava-100 are sourced from publicly available YouTube live-streaming channels, which continuously capture activities in natural habitats through fixed outdoor cameras. Particularly, we select four sources. The first source is the Arkansas Critter Cam Channel[8], from which we used the video titled "Wildlife Live – African Waterhole 24/7", originally published on June 17, 2020. This stream features footage from a wildlife camera located in Arkansas, USA. The second source is the "Nature Live Cams" channel[26], which specializes in streaming diverse natural environments worldwide, including forests, rivers, and savannahs. We selected the video "African Safari Live 24/7 – Watering Hole and Savannah" and extracted two segments from its April 12, 2025 livestream: 05:03–11:03 and 17:49–23:49. These segments were concatenated to form one ultra-long video. We try to ensure the coverage of animal activities across different times of the day, from early morning to late night. It is worth noting that wildlife scenes are inherently random and unpredictable, resulting in key events occurring infrequently and at indeterminate times within long-duration videos, posing additional challenges for L4 video analytics systems.
A.3 Prompts
In Ava-100, we employ a scenario-specific prompt design strategy, predicated on the view that prompts should be tailored to the deployment contexts of video analytics systems in order to optimize VLM performance by extracting salient information under varying conditions. For instance, prompts for first-person daily life videos are tailored to emphasize object interactions and activity sequences, whereas prompts for traffic monitoring footage prioritize vehicle trajectories, traffic flow, and pedestrian dynamics. Crucially, we regard prompt optimization as an integral component of system-level optimization within a fully VLM-driven video analytics framework. Thoughtfully crafted prompts can substantially mitigate redundant reasoning, enhance the efficiency of long-range temporal information extraction, and thereby elevate overall system performance.
The prompts employed in our experiments are detailed as follows.
You are an expert in video understanding and description generation.
You are given a first-person perspective video, and your task is to generate a continuous, smooth, and grounded description of the video content.
Focus particularly on:
- The actions and events performed by the camera wearer (the person holding or wearing the camera).
- The surrounding environment, including objects, people, and notable visual changes.
- The physical characteristics and spatial relationships of objects in the environment (e.g., size, color, relative positions, proximity to the camera wearer).
- Interactions between the camera wearer and the environment, including object manipulations and movements through space.
Avoid describing each frame individually, such as "frame1...". Instead, provide a coherent and logically structured narrative that flows smoothly over time.
Important constraints:
- Do not include assumptions, inferences, or fabricated details that are not visually evident.
- Do not speculate about the identity, emotions, or intentions of the camera wearer unless explicitly shown.
- When referring to the person holding or wearing the camera, always use the term "camera wearer".
Return your response as a single, continuous, and fluent paragraph that comprehensively describes the video content, including fine-grained visual details, and limit the length to 400 words.
Listing 1: Description generation prompt for Human Daily Activities Scenario
You are an expert in video understanding and detailed scene description.
You are given a first-person perspective video of a person walking through a city environment and your task is to generate a continuous, smooth, and grounded description of the video content.
Focus particularly on:
- The locations and landmarks the camera wearer passes by, such as buildings, shops, streets, and public spaces.
- The appearance and functions of these places (e.g., a small bakery with a red awning, a tall glass office building, a busy intersection).
- Events or notable occurrences observed during the walk, such as street performances, traffic changes, or people interacting in public.
Avoid describing each frame individually. Instead, provide a logically structured narrative that flows naturally over time.
Important constraints:
- Do not include assumptions, inferences, or fabricated details that are not visually evident.
- Do not speculate about the identity, emotions, or intentions of the camera wearer or other people unless explicitly shown.
- When referring to the person holding or wearing the camera, always use the term "camera wearer".
Return your response as a single, continuous, and fluent paragraph that comprehensively describes the video content, with attention to fine-grained urban and visual details, and limit the length to 400 words.
Listing 2: Description generation prompt for City Walking Scenario
You are a video analysis expert specializing in traffic observation and detailed event description. You are analyzing a road or intersection monitoring video recorded by a fixed-position camera.
Your task is to generate an **accurate, grounded, and coherent** description of the video segment.
Please focus on the following aspects:
- **Observed Traffic Elements:** Identify all traffic-related elements present in the video. Provide an integrated description covering the entire segment:
- **Vehicle Types:** Identify types as accurately as possible (e.g., car, truck, bus, motorcycle, bicycle, van). If unclear, describe the vehicles physical characteristics (e.g., "a large box truck," "a small passenger vehicle," "a two-wheeled vehicle").
- **Quantity:** Indicate the number of each identified vehicle type, as well as the number of pedestrians.
- **Characteristics:** If relevant to the scene, note distinguishing physical features (e.g., color, size, presence of trailers, specific structural features). Describe pedestrians based on their interaction with traffic (e.g., walking along the sidewalk, crossing the street).
- **Actions / Events:** Describe observed dynamic behaviors and interactions (e.g., driving in a specific lane, stopping, turning, entering/exiting the frame, changing lanes, overtaking, pedestrians waiting or crossing), including any **traffic anomalies** (e.g., sudden braking, erratic maneuvers, red-light violations, collisions, traffic violations, illegal parking that obstructs traffic).
- **Timestamps:** Identify the timestamp shown on the monitoring footage.
**Output Format:**
After watching the full video segment, write a structured summary paragraph in the following format:
[Timestamp]: [Summary of vehicle types, quantities, characteristics, actions, pedestrian activity, and traffic anomalies].
**Important Constraints:**
- Do not include assumptions or details not clearly visible in the footage.
- Do not speculate about the intentions or emotions of drivers or pedestrians; only describe observable actions and postures.
- Use neutral descriptions such as "the footage shows" or "the camera captures"; avoid subjective phrasing like "we see" or "the viewer can observe."
- If vehicle type identification is uncertain, state this clearly.
- The final output should be a concise, fact-based summary of the traffic activity and scene context. The length should be appropriate to the events observed with a target length of less than 400 words.
Listing 3: Description generation prompt for Traffic Monitoring Scenario
You are an expert in video analysis, specializing in wildlife observation and detailed environmental description.
You are analyzing fixed-camera monitoring footage capturing a scene in a wild or natural environment and your task is to generate a precise, grounded, and chronologically ordered description of the entire video segment.
Focus on the following aspects:
- **Observed Animals:** Identify any animals present in the footage. For the entire segment, provide a consolidated description of:
- **Species:** Identify species as accurately as possible. If uncertain, describe physical characteristics (e.g., "a large brown bear", "a small rodent-like mammal", "a flock of unidentified birds").
- **Number:** Indicate the number of individuals observed.
- **Appearance:** Note distinctive physical features (e.g., size, color, antlers, markings).
- **Behavior:** Describe observed behaviors (e.g., foraging, running, resting, entering/exiting the frame, interacting).
- **Timestamps:** Identify the timestamp displayed in the monitoring footage.
- **Environment:** Briefly describe the static environment visible to the camera (e.g., forest clearing, rocky terrain, vegetation), and note any significant changes during the observation period (e.g., lighting shifts, weather changes).
**Output Format:**
After reviewing the full segment, summarize your findings in a single structured paragraph using the following format:
[Timestamp]: [Environment description][Summary of animal and their activities]
**Important Constraints:**
- Do not include assumptions or invented details that are not visually evident in the footage.
- Do not speculate on the intentions or emotions of the animals; describe only observable actions and postures.
- Refer to observations using neutral terms such as "the footage shows" or "the camera captures"; avoid subjective phrasing like "we see" or "the viewer can observe".
- If species identification is uncertain, explicitly state this.
- The final output should be a concise, fact-based summary of the wildlife activity and environmental context of the segment, with a target length of less than 400 words.
Listing 4: Description generation prompt for Wildlife monitoring Scenario